OpenStudy (anonymous):

Please help!!:/ Divide: p^2-q^2/p+q divided by p-q/p+q

OpenStudy (mertsj):

\[\frac{p^2-q^2}{p+q} \times \frac{p+q}{p-q}\]

OpenStudy (mertsj):

Factor the first numerator and cancel the common factors.

OpenStudy (anonymous):

Thank you so much!!:)

OpenStudy (mertsj):

yw